For each of the following numbers, find the smallest whole number by which it should be multiplied so as to get a perfect square number. Also find the square root of the square number so obtained. 
1458

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student
1458 = 2×3×3×3×3×3×3
= (3×3)×(3×3)×(3×3)×2
Here, 2 cannot be paired.
∴We will multiply 1458 by 2 to get perfect square. New number = 1458 × 2 = 2916

2916 = 2×2×3×3×3×3×3×3
⇒2916 = (3×3)×(3×3)×(3×3)×(2×2)⇒2916 = 32×32×32×22
⇒2916 = (3×3×3×2)2
⇒ √2916 = 3×3×3×2 = 54
Thanks